3. Count von Count likes to (loudly) count passing cars next to Oscars home. This creates a negative externality via noise pollution that irritates Oscar. The marginal private cost of counting is 40, and marginal damage is given by 2Q, so: MPC 40 MD 2Q a) Explain why the existence of externalities can be considered a market failure. b) lgnoring the externality, find the privately optimal quantity (Qo) of counting if the marginal benefit is given by: MB 60 2Q. Assume Q is in terms of hours a week. c) What is the equation for marginal social cost (MSC)? d) Find the socially optimal quantity of hours of car counting (Q*)
Answer #1

(a) In presence of positive (negative) externality, the firms produce lower than (higher than) the socially efficient output at lower than (lower) than the socially efficient price. Therefore, when there is positive (negative) externality, a free-market equilibrium leads to underproduction (overproduction), which is a misallocation of resources, leading to market failure.

(b) Equating MB with MPC,

60 - 2Q = 40

2Q = 20

Q = 10 (Q0)

(c) MSC = MPC + MD

MSC = 40 + 2Q

(d) Equating MB with MSC,

60 - 2Q = 40 + 2Q

4Q = 20

Q = 5 (Q*)